# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ Appendix I Models We Use in Experiments
0.610
0.719
(±0.008)
0.629
0.744
(±0.006)
0.612 0.737
Qwen Qwen2.5-7B acc 0.504
0.622
(±0.010)
0.499
0.623
(±0.009)
0.554
0.635
(±0.007)
0.519 0.626
acc_
norm
0.527
0.637
(±0.011)
0.521
0.633
(±0.012)
0.576
0.650
(±0.005)
0.541 0.640
Qwen2.5-7B- Instruct acc 0.568
0.713
(±0.007)
0.545
0.700
(±0.004)
0.543
0.707
(±0.005)
0.552 0.707
acc_
norm
0.587
0.727
(±0.006)
0.576
0.713
(±0.002)
0.565
0.721
(±0.007)
0.576 0.720
Mistral Mistral-7B- v0.3 acc 0.463
0.640
(±0.004)
0.479
0.636
(±0.007)
0.443
0.670
(±0.004)
0.462 0.649
acc_
norm
0.487
0.654
(±0.003)
0.499
0.650
(±0.008)
0.473
0.685
(±0.006)
0.486 0.663
Mistral-7B- Instruct-v0.3 acc 0.566
0.767
(±0.006)
0.667
0.754
(±0.008)
0.606
0.785
(±0.004)
0.613 0.768
acc_
norm
0.593
0.779
(±0.006)
0.686
0.767
(±0.009)
0.631
0.796
(±0.006)
0.636 0.781
average acc 0.511 0.686 0.539 0.672 0.529 0.689 0.526 0.682
acc_
norm
0.565 0.700 0.586 0.685 0.578 0.704 0.576 0.696
Table 14: Total zero-shot and few-shot evaluation results on Thunder-NUBench. SD denotes standard deviation across random seeds or runs. acc denotes the accuracy, and acc_norm indicates the length-normalized accuracy. In all few-shot experiments, results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and five examples of validation set data (5 shots).
Appendix M Total Model Performance on Thunder-NUBench
This section reports zero-shot, 5-shot performance (averaged over five random seeds: 42, 1234, 3000, 5000, and 7000) of various pre-trained language models on Thunder-NUBench, prior to any fine-tuning. The default prompt is prompt 1, the instruction prompt is prompt 7, and the reasoning prompt is prompt 11 from Table 10. Table 14 shows the overall results.
Instruction-tuned models consistently outperform their pretrained counterparts across all model series. Although detailed instruction prompts provide explicit task descriptions, the substantial performance improvement observed with few-shot compared to zero-shot settings suggests that concrete examples are even more effective in helping models understand and perform the task. This trend holds even under instruction-based prompting, indicating that examples play a more critical role than instructions alone in guiding model behavior on logical negation tasks. We can also observe that performance varies substantially depending on the prompt type. For example, Mistral-7B-Instruct-v0.3 achieves an accuracy of 0.566 with the default prompt in a zero-shot setting, but this rises to 0.667 with the instruction prompt and 0.606 with the reasoning prompt. Notably, both results on instruction and reasoning prompts significantly outperform the ones on default prompt. This pattern indicates that each model has different preferences for prompt type, and selecting the optimal prompt is crucial for maximizing performance.
Appendix N Total Model Performance after SFT
zeroshot
fewshot
(±SD)
zeroshot
sft
3B-4B Llama Llama-3.2-3B acc 0.463
0.616
(±0.005)
0.792
acc_norm 0.483
0.633
(±0.004)
0.800
Llama-3.2-3B-Instruct acc 0.490
0.619
(±0.006)
0.808
acc_norm 0.512
0.633
(±0.001)
0.815
Qwen Qwen2.5-3B acc 0.495
0.600
(±0.009)
0.747
acc_norm 0.526
0.622
(±0.009)
0.753
Qwen2.5-3B-Instruct acc 0.667
0.758
(±0.006)
0.765
acc_norm 0.686
0.771
(±0.006)
0.775
average acc 0.529 0.648 0.789
acc_norm 0.552 0.665 0.778
zeroshot
fewshot
(±SD)
zeroshot
sft
7B-8B Llama Llama-3.1-8B acc 0.477
0.643
(±0.005)
0.851
acc_norm 0.504
0.657
(±0.008)
0.860
Llama3.1-8B-Instruct acc 0.564
0.698
(±0.007)
0.855
acc_norm 0.584
0.706
(±0.006)
0.868
gemma gemma-7b acc 0.492
0.621
(±0.005)
0.847
acc_norm 0.518
0.635
(±0.006)
0.849
gemma-7b-it acc 0.588
0.700
(±0.007)
0.838
acc_norm 0.610
0.719
(±0.008)
0.845
Qwen Qwen2.5-7B acc 0.499
0.623
(±0.009)
0.820
acc_norm 0.521
0.633
(±0.012)
0.825
Qwen2.5-7B-Instruct acc 0.545
0.700
(±0.004)
0.783
acc_norm 0.576
0.713
(±0.002)
0.784
Mistral Mistral-7B-v0.3 acc 0.479
0.636
(±0.007)
0.839
acc_norm 0.499
0.650
(±0.008)
0.851
Mistral-7B-Instruct-v0.3 acc 0.667
0.754
(±0.008)
0.815
acc_norm 0.686
0.767
(±0.009)
0.823
average acc 0.539 0.672 0.831
acc_norm 0.586 0.685 0.838
Table 15: Comparison of negation understanding performance on Thunder-NUBench across 3B–4B and 7B–8B models. Each model is evaluated in zero-shot, few-shot (5-shot average with standard deviation), and post-supervised fine-tuning (SFT) zero-shot settings. Metrics include raw accuracy (acc) and normalized accuracy (acc_norm). All evaluations use the instruction-based prompt format, prompt 7 from Table 10, consistent with the SFT data format.
We present model performance on Thunder-NUBench after supervised fine-tuning, highlighting improvements from task-specific training. Because SFT is performed on instruction-based data, only the instruction-based prompt format (prompt 7 from Table 10) is used for evaluation. We assess zero-shot performance exclusively to directly measure the model’s ability from fine-tuning without the influence of in-context examples.
Note that we exclude both gemma-3-4b-it and gemma-3-4b-pt from our SFT experiments. During preliminary fine-tuning using LoRA, these models consistently produced unexpected errors across all configurations. Such failures are not limited to a specific checkpoint or setting, but appear to be a universal issue affecting the entire Gemma-3 series (Team et al., 2025). These errors stem from a structural incompatibility between the model’s uploaded configuration and the expectations of the transformers library’s auto-loading mechanism.  [^7] https://github.com/huggingface/transformers/issues/36683. We have initially selected these models to explore performance within the 3-4B parameter range in the Gemma family, but due to the unresolved errors during SFT, we are unable to include them in our final evaluation. Future work may revisit these models once the associated issues are addressed.
Table 15 presents model performance on Thunder-NUBench across three evaluation settings: zero-shot, few-shot (with 5-shot averaging), and after supervised fine-tuning (SFT). Supervised fine-tuning leads to substantial performance improvements compared to both baseline zero-shot and few-shot conditions. Across nearly all models and sizes, the SFT-zero-shot setting consistently outperforms few-shot baselines.
Appendix O ScoNe Performance after Thunder-NUBench Fine-Tuning
model
size
model
series
Model Name
Baseline
zeroshot
SFT
(on Thunder-NUBench)
zeroshot
3-4B Llama Llama-3.2-3B 0.585 0.546
Llama-3.2-3B-Instruct 0.577 0.676
Qwen Qwen2.5-3B 0.765 0.760
Qwen2.5-3B-Instruct 0.744 0.792
3-4B models average 0.668 0.694
7-8B Llama Llama-3.1-8B 0.620 0.544
Llama-3.1-8B-Instruct 0.696 0.759
gemma gemma-7b 0.586 0.601
gemma-7b-it 0.740 0.743
Qwen Qwen2.5-7B 0.758 0.769
Qwen2.5-7B-Instruct 0.778 0.794
Mistral Mistral-7B-v0.3 0.544 0.502
Mistral-7B-Instruct-v0.3 0.740 0.740
7-8B models average 0.683 0.682
total average 0.678 0.686
Table 16: Zero-shot performance on the ScoNe benchmark before and after supervised fine-tuning (SFT) on the Thunder-NUBench training set. Each model is evaluated without seeing ScoNe data during training.
ScoNe (Scoped Negation Dataset) is a benchmark designed to evaluate whether models truly understand the semantic scope of negation in natural language, using controlled contrast sets in a natural language inference (NLI) framework with up to two negation morphemes per example. ScoNe consists of a train and test set, each of which has 5,010 and 100 instances (She et al., 2023).
In this experiment, we first measure the zero-shot baseline performance of various models on the ScoNe dataset. We then evaluate the same models after supervised fine-tuning (SFT) on the Thunder-NUBench training set, without ever exposing them to ScoNe during training. This setup allows us to assess whether negation understanding learned from Thunder-NUBench generalizes to ScoNe.
Overall, we observe that performance improves moderately after fine-tuning (total average of accuracy: 0.678 $\rightarrow$ 0.686). This suggests that negation understanding learned from Thunder-NUBench transfers effectively to a structurally different evaluation set like ScoNe. Instruction-tuned models benefit the most from SFT on Thunder-NUBench, likely because the fine-tuning data follows an instruction format, to which they are already well adapted. On the other hand, some pretrained models exhibit performance degradation.
Appendix P Total Analysis of Model Predictions on the Thunder-NUBench
We provide a detailed analysis of model predictions across negation types and sentence structures, examining common error patterns and behavioral trends.
Table 17 presents the incorrect choice distribution and confusion rates for local negation types across a range of 3-4B and 7-8B pretrained and instruction-tuned models under zero-shot and few-shot settings. We report few-shot results using a fixed random seed (1234), which corresponds to the default seed used in the LM Evaluation Harness framework. Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make fine-grained confusion analysis less interpretable.
A notable observation is that choice4 (paraphrase) was rarely selected as the incorrect answer, suggesting that models can generally distinguish paraphrases from negation. In particular, the Qwen2.5-7B-Instruct model under a few-shot setting achieved a remarkable result, never selecting choice4 as the incorrect answer. This implies that even when structural transformations or synonyms are introduced, experimented models do not confuse them with negation, possibly because paraphrases tend to preserve the positive polarity of the original sentence, in contrast to the strong semantic shift introduced by local negation or contradiction.
Across all models, the most frequent source of confusion was choice2 (local negation), often selected as the answer instead of standard negation. This suggests that models struggle to differentiate sentence-level negation from subclausal negation. Within local negation types, the compound_part category exhibited the highest confusion rates across many models (e.g., 58.12% for Mistral-7B-v0.3 zero-shot, 56.68% for Llama-3.2-3B zero-shot), indicating that when negation applies to one clause within a compound sentence, it is particularly challenging for models to resolve its scope relative to full-sentence negation.
Incorrect Choice Distribution choice2 confusion rate
Error rate

### v3

@@ 8 Paraphrased sentence:
0.689
(±0.013)
2-3B models
average
0.430 0.377 0.453 0.352 0.545 0.299 0.582 0.332 0.584 0.554
7-8B models
average
0.499 0.468 0.524 0.489 0.605 0.534 0.643 0.560 0.685 0.645
api models
average
- 0.686 - 0.652 - 0.687 - 0.743 - -
base models
average
0.438 0.376 0.462 0.397 0.544 0.440 0.585 0.447 0.637 0.599
instruct-tuned
models average
0.501 0.483 0.525 0.464 0.614 0.427 0.648 0.477 0.646 0.613
total average 0.469 0.474 0.493 0.469 0.579 0.478 0.617 0.512 0.642 0.606
Table 13: Zero-shot, few-shot, and SFT evaluation results on Thunder-NUBench with the detail prompt. SD denotes standard deviation across random seeds or runs. Few-shot results are averaged over five random seeds (42, 1234, 3000, 5000, and 7000) and one, five, and ten demonstration examples (1-, 5-, 10-shot). Red text indicates the model with the highest performance in each column (excluding API models).
Following the classification of prompts introduced in Appendix O, we report results on Thunder-NUBench separately for the simple, definition, and detail prompt styles. The experiments are conducted under three evaluation regimes: (i) zero-shot, (ii) few-shot (1-, 5-, and 10-shot, averaged across seeds), and (iii) supervised fine-tuning (SFT). For SFT, we assess models only in the zero-shot setting in order to directly measure the effect of task-specific training without the influence of in-context examples.
Figure 4: Model performance on Thunder-NUBench with definition prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Figure 5: Model performance on Thunder-NUBench with detail prompt. Circles (blue) represent the average performance of 2-3B models, squares (purple) indicate the average for 7-8B models, upward triangles (orange) signify the average of base models, and downward triangles (red) denote the average of instruction-tuned models. Stars (green) represent API models.
Table 11, Table 12, and Table 13 present the complete numerical results for the simple, definition, and detail prompts, respectively. The same results are also summarized in graphical form in Figure 3, Figure 4, and Figure 5.
Across prompt types, we observe the following trends:
  • Simple prompts. Results for simple prompts are described in Section 5.2; we summarize them here only for completeness.
  • Definition prompts. Performance patterns closely resemble those of simple prompts. However, in the option-selection evaluation, 2-3B models show smaller gains from additional shots, while 7-8B models tend to achieve slightly higher accuracy after SFT compared to 10-shot prompting.
  • Detail prompts. For smaller (2-3B) models in the option-selection evaluation, adding more shots can actually reduce performance, reflecting difficulties with long and complex prompt instructions to these small models. Nonetheless, these models, along with larger ones, show larger SFT gains under detail prompts than under simple or definition prompts.
Overall, these findings show that prompt formulation plays a non-trivial role in shaping performance trends. In all cases, SFT consistently boosts zero-shot performance.
Appendix Q General Benchmark Performance after SFT
tasks
arc_
challenge
arc_
easy
gsm8k hellaswag
openbook-
qa
wino-
grande
avg
metric acc acc
exact_
match
acc_norm acc_norm acc
Before SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.432 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.756 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.702 0.607
After SFT with simple prompt on Thunder-NUBench
2-3B models
average
0.431 0.746 0.365 0.713 0.411 0.671 0.556
7-8B models
average
0.524 0.819 0.536 0.798 0.453 0.721 0.642
base models
average
0.474 0.788 0.428 0.772 0.437 0.706 0.601
instruct-tuned
models average
0.495 0.787 0.497 0.751 0.432 0.693 0.609
total average 0.484 0.788 0.463 0.762 0.435 0.699 0.605
Before SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with definition prompt on Thunder-NUBench
2-3B models
average
0.433 0.745 0.341 0.713 0.407 0.669 0.551
7-8B models
average
0.528 0.819 0.545 0.799 0.455 0.725 0.645
base models
average
0.477 0.785 0.419 0.772 0.436 0.709 0.600
instruct-tuned
models average
0.498 0.790 0.496 0.753 0.433 0.694 0.611
total average 0.487 0.788 0.457 0.762 0.434 0.701 0.605
Before SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.431 0.747 0.325 0.714 0.409 0.668 0.549
7-8B models
average
0.514 0.810 0.592 0.797 0.458 0.727 0.65
base models
average
0.464 0.784 0.480 0.768 0.437 0.709 0.607
instruct-tuned
models average
0.493 0.782 0.475 0.755 0.437 0.694 0.606
total average 0.479 0.783 0.478 0.762 0.437 0.701 0.607
After SFT with detail prompt on Thunder-NUBench
2-3B models
average
0.429 0.746 0.359 0.713 0.409 0.668 0.554
7-8B models
average
0.523 0.816 0.551 0.8 0.458 0.723 0.645
base models
average
0.469 0.785 0.424 0.773 0.439 0.704 0.599
instruct-tuned
models average
0.496 0.787 0.513 0.753 0.434 0.695 0.613
total average 0.483 0.786 0.469 0.763 0.437 0.699 0.606
Table 14: Average performance of 2–3B / 7–8B, and base / instruction-tuned models on six general benchmarks before and after SFT on Thunder-NUBench. Here, acc denotes accuracy, acc_norm is normalized accuracy, and exact_match requires an exact string match with the reference answer.
To assess whether supervised fine-tuning (SFT) on Thunder-NUBench affects performance on broader natural language understanding tasks, we evaluate models on six widely used benchmarks: ARC-Challenge, ARC-Easy, GSM8K, HellaSwag, OpenBookQA, and WinoGrande (clark2018think; cobbe2021training; zellers2019hellaswag; mihaylov2018can; sakaguchi2021winogrande). These datasets cover a diverse range of domains, including commonsense reasoning, scientific knowledge, mathematics, and reading comprehension.
Following the classification of prompts introduced in Appendix O, SFT is conducted separately for the simple, definition, and detail prompt styles. We then report post-SFT zero-shot performance across the six benchmarks to verify whether task-specific training on Thunder-NUBench preserves general capabilities.
Table 14 summarizes the results. We find that performance on general benchmarks remains broadly stable after SFT, indicating that fine-tuning on Thunder-NUBench does not substantially harm general capabilities.
Appendix R Total Analysis of Model Predictions on the Thunder-NUBench
This appendix extends the error analysis presented in Section 5.3, providing the complete results. In particular, we examine (i) incorrect choice distributions and (ii) confusion rates for local negation categories, comparing models of different sizes (2-3B vs. 7-8B) and training paradigms (pretrained vs. instruction-tuned) under both zero-shot and few-shot conditions.
We report few-shot results using a fixed random seed (1234), which corresponds to the default seed used in the LM Evaluation Harness framework. Averaging over multiple seeds was avoided, as it could obscure specific error patterns and make fine-grained confusion analysis less interpretable.
For consistency with the main text, we report results only under the simple prompt, which serves as the default evaluation setting throughout the paper.
We organize the complete prediction analysis by model family.
  • Results for the Gemma family are reported in Table 15, Table 16.
  • Results for the LLaMA family are reported in Table 17, Table 18.
  • Results for the Mistral family are reported in Table 19, Table 20.
  • Results for the Qwen family are reported in Table 21, Table 22.
  • Results for API models are reported in Table 23.
Each table follows the same format, reporting error rates, incorrect choice distributions, and confusion rates across local negation subcategories under zero-shot, few-shot, and SFT conditions.
In the option-selection evaluation setting, we also track cases labeled as “Answer Format Wrong.” This category captures instances where the model’s output does not follow the required answer format (selecting strictly one of A, B, C, or D). Because such responses cannot be mapped to a specific incorrect option, they are not included in the incorrect choice distribution or confusion rate. Instead, we report their raw counts alongside the other error statistics. This also serves as an indicator of the model’s ability to follow output-format instructions.
Overall, we observe consistent trends that larger models (7-8B) achieve lower error rates than smaller ones (2-3B), and introduction-tuned variants generally outperform base models. Supervised fine-tuning (SFT) tends to reduce error rates. Meanwhile, incorrect choice distributions and local negation confusion rates vary across models and settings, showing that this analysis can serve as a useful tool to identify which aspects of negation remain particularly challenging.
Notably, we also find some unusual behaviors. For example, the Gemma-1.1-2B-it model after SFT produces 1,239 outputs with incorrect answer formats in the option-selection setting, indicating formatting issues rather than genuine reasoning errors. Similarly, for API models, Claude 3.5 Haiku shows increasing formatting errors as more shots are added, which directly degrades performance.
Error
Rate
Incorrect Choice Distribution Local Negation Confusion Rate